OpenStudy (anonymous):

in how many ways can 3 boooks be selected from a shelf of 25 books?

OpenStudy (anonymous):

click good answer

OpenStudy (anonymous):

2300 ways

OpenStudy (campbell_st):

25C3

OpenStudy (anonymous):

25!/3!(25-3)!=25!/3!22!=25*24*23/3!=25*24*23/6=25*4*23=100*23=2300